OpenStudy (anonymous):

find the length of a diagonal of a square with sides 10in long

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 10^2 + 10^2 = c^2 100 + 100 = c^2 200 = c^2 c^2 = 200 c = sqrt(200) c = sqrt(100*2) c = sqrt(100)*sqrt(2) c = 10*sqrt(2) So the diagonal is 10*sqrt(2) inches long.